You are the owner of a car dealership and you practice “no haggle” sales policy. Last year, you made a record profit of GHC 1.5 million. Your dealership competes in a market with price elasticity of demand of -1.3. Your marginal cost is GHC 12000. How much should you charge to maintain your record profit?


1
Expert's answer
2021-09-27T09:06:30-0400

To find record profit we will apply formula

 "P = \frac{EF}{ 1 + EF}\\

 

where\\ EF = \frac{NEM}{ (1 + NEM}"

Where =EF = Elasticity for a firm ,  

 EM = Elasticity of whole market,     

N=number of players in market

 MC is Marginal cost of firm and P is Price

 

There are other two dealers in the market as well. Hence N = 3. Here ed = -1.3. This indicates that we have

 "P = MC \times \frac{Ned }{ (Ned + 1)}\\

 

= 12000 x\times \frac{3\times (-1.3) }{ (3\times (-1.3) + 1))}\\

 

= GHC16,137.93"

Thus, the price you should charge for a maintain automobile is GHC16,137.93